### v3

@@ Abstract
The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns. Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users. Such explanations are essential in enhancing trustworthiness and safety. Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions. This work presents an accountability and explainability architecture implemented for mobile Robot Operating System (ROS)-based robots. The proposed solution consists of two main components. First, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology. Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box. The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities. This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing challenges inherent in the use of autonomous agents in real-world scenarios.
keywords:
@@
{highlights}
Security improvements for autonomous agents in public spaces.
Review of blockchain for immutable logs, EDRs, XAI and blockchain, XARs, and LLMs.
Accountability and Explainability architecture for autonomous agents.
Comprehensive assessment across diverse navigation scenarios and metrics.
Future research directions for accountable and explainable autonomous agents.
1 Introduction
The rising presence of mobile robots alongside humans has increased concerns about security and trust. One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios where their behavior may impact human well-being [1]. Transparency and reliability in these justifications are essential [2], as such properties enable effective human-robot interaction and foster confidence, particularly for non-expert users who require clear and understandable explanations [3, 4]. However, the complexity of robotic decision-making and the need for thorough post-event analysis make achieving these goals particularly difficult.
To address these challenges, accountability must be ensured through a robust, tamper-proof logging system. The most substantial evidence for post-event analysis comes from logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect data tampering or unauthorized breaches. These systems must guarantee the integrity, confidentiality, and availability of data while enabling incident detection, compliance, and forensic investigations [5].
However, covering tracks in log systems has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions. As a result, it is necessary to strengthen system properties against unauthorized modifications [6]. This includes safeguarding the integrity of log data and implementing measures to detect and counteract attempts at tampering. Effective accountability solutions must ensure the detection and isolation of faulty behaviors and their origin. Moreover, they should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
Integrating accountability solutions into robotic systems often demands a trade-off with system performance. Specifically, computational power and networking bandwidth may be adversely affected [8]. The inclusion of anti-tampering techniques, such as full disk and communication encryption, can be restrictive due to resource limitations [9]. Additionally, high-cost tamper-proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
@@ 1 Introduction
To this end, blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability. The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction or manipulation of data through illicit channels. The nature of this technology mitigates tampering risk, safeguarding data [11].
@@
Nevertheless, while blockchain secures log data, the evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events. These factors may result in limited usefulness and misalignment with the expectations of non-expert users [12]. The raw log messages produced by autonomous agents are usually noisy and semi-structured in format. In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages [13]. Addressing this issue requires methods to organize and filter log data, making it accessible and meaningful.
@@
In this context, explaining the reasons behind a robot’s specific and unexpected actions is essential for diagnosing failures and comprehending deviations from intended goals. Equally important is ensuring these explanations are understandable to users [14]. Explainable Artificial Intelligence (XAI) is a research area dedicated to enhancing the transparency and interpretability of Artificial Intelligence (AI) systems to ensure their responsible and ethical deployment. As autonomous robots grow in complexity, understanding and predicting their behavior has become increasingly challenging for non-expert users. This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents [15]. XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity and user comprehension [16, 17].
@@
To further improve the effectiveness of XAR systems, Large Language Models (LLMs) excel in diverse natural language processing tasks, including translation, coding, or text summarization [18, 19, 20]. Our approach integrates LLMs with blockchain-secured log data. By using retrieval-based techniques, this solution generates accurate and context-aware natural language explanations based on verifiable information [21]. This integration ensures that explanations are clear, coherent, and grounded in a traceable and reliable context, effectively addressing user comprehension and data integrity concerns.
This work presents an accountability and explainability architecture implemented for Robot Operating System (ROS)-based mobile robots. This solution includes two main components. First, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component. The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events. Both components are connected through an additional module responsible for processing raw data collected by the first one, eliminating non-essential information, and building connections and relations between data. These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations for the code generated and the behaviors developed by the autonomous agent. To validate our approach, we conducted a comprehensive evaluation, assessing the architecture’s impact on system performance and analyzing the quality of the generated explanations. The evaluation considered key metrics, including correctness, accuracy, helpfulness, coherence, and user satisfaction. Our results demonstrate that the architecture effectively balances accountability and explainability without compromising the robot’s performance, providing a scalable and reliable framework for real-world applications.
The remainder of this paper is organized as follows: Section 2 provides an overview of previous works and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots. Section 3 delves into the system architecture, offering a detailed description of the designed approaches. Section 4 details the technical elements of the designed system attending to the proposed architecture. Section 5 discusses the experimental process, the scenarios and the variables employed for assessment. Section 6 outlines the obtained results, particularly addressing accountability and explainability metrics. Finally, Section 7 summarises the obtained conclusions.
2 Related Works
In this section, we explore key works related to various domains relevant to our approach. This includes advances in how blockchain ensures the immutability of logs, research on applying Event Data Recorders (EDRs) in robotics, significant contributions to integrating XAI with blockchain, and XAR. Finally, we examine the evolution of LLMs for generating natural language explanations and their advancements within robotics.
2.1 Blockchain for Immutable Logs
Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks. These blocks are linked through hashing and the storage of previous block content, ensuring the integrity of transactions. Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process. As a result, data stored in the blockchain is immutable [22, 23].
There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility. Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification [24]. While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible, and scalable consensus protocol [25, 26].
Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or InterPlanetary File Systems (IPFSs), which may require central authorities for storage and verification [27, 28]. Blockchain offers consensus and immutability, coupled with the capability of integrating logic mechanisms such as Smart Contracts (SCs). SCs contain functions that either write values onto the blockchain, thus altering its state, or read data from it [29]. These features make blockchain the preferred choice for applications where data integrity and trust are required [30].
Different secure logging systems rely on permissioned blockchain frameworks for log immutability [31], introducing accountability even in embedded systems [7]. To get a balance between throughput and cost-effectiveness, some works combine both categories [32], developing approaches that include a permissioned blockchain to ensure per-entry immutability and public non-repudiation solution [33]. Nevertheless, inherent features of permissionless blockchain such as transparency and trust, have also boosted the development of secure logging systems [34].
Our architecture uses the immutability of blockchain to store evidence of the integrity of recorded data, thus providing real-time guarantees against manipulation and ensuring that the explanations generated by our system are always based on non-manipulated data.
2.2 Blockchain for Event Data Recorders in Robotics
EDRs, or black boxes, are essential for capturing and analyzing events in robotic systems. These recorders help to comprehend system behavior, detect anomalies, and ensure safety [35, 36]. Therefore, these solutions are needed in forensic investigations and ethical assessments regarding autonomous agents [37]. The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [38, 9]. These solutions typically include functionalities such as hashing, signing, and encryption [39], while ensuring event integrity through anti-tamper mechanisms [40].
Using blockchain technology in EDRs ensures the integrity of a transparent and auditable record of events. These features enhance accountability and trust in the recorded data by preventing tampering [41]. Decentralized and distributed storage further eliminates reliance on central authorities and mitigates data loss risks.
Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management. The Black Block Recorder (BBR) system uses permissioned blockchain for immutable logging, combining Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity [42]. Blockchain applications in robot swarms, including secure communication, data logging, and consensus agreements, are discussed in [43]. In [44], a time-segmented consortium blockchain for robotic event registration is proposed, enabling efficient event organization and retrieval within specific time frames. Authors in [45] present a tool to bridge ROS and the Ethereum blockchain, evaluating encryption efficiency and stability, transaction response time, and packet loss rate. Nevertheless, these works lack the capability to save integrity proofs in the blockchain at selected intervals, enhance granularity, or include service definitions or SCs to improve the verifiability process.
2.3 Explainable Artificial Intelligence and Blockchain
XAI aims to align AI decisions with human cognitive expectations by employing techniques such as contextual relationships, contrastive explanations, and methods like counterfactual reasoning and interpretable models that enhance interpretability and transparency [46, 47]. Insights from social sciences show people tend to interpret explanations more positively when they reflect human-like reasoning, emphasizing intentions, goals, and contextual factors rather than purely technical or statistical descriptions [48]. By combining XAI with blockchain transparent record-keeping, explanations can be retraced and validated, ensuring both interpretability and accountability in AI-driven systems [49, 50].
The benefits of this integration are remarkable in various domains. In healthcare, XAI and blockchain have been combined for secure, transparent patient data management [51], logical disease prediction, and immutable medical diagnostics records [52]. Similarly, in the legal domain applications, blockchain and XAI support tamper-proof decision auditing, using IPFS and Ethereum for cryptographic proofs that enable data verification and accountability [53, 54, 28]. Privacy-preserving systems with homomorphic encryption and semantic ontologies further ensure data protection while maintaining explainability [55].
Regarding autonomous agents like unmanned aerial vehicles or autonomous vehicles, blockchain and XAI ensure secure and transparent decision-making [56]. This integration enhances collaboration, even in complex and time-critical scenarios [57]. Decentralized edge models for collaborative AI, leveraging blockchain for secure communication and real-time data sharing, further optimize resources and enhance reliability in these environments [58].
Our work introduces Rosbag2 files as an alternative to IPFS or distributed file systems for storing contextual data in autonomous robotic systems. This approach supports reproducibility and enables a more precise analysis of robotic behaviors [42]. Rosbag files provide a standardized format for recording large volumes of robotic data, facilitating comprehensive interaction capture for improved analysis. To address the indeterminism of LLM algorithms, we propose using a permissionless blockchain to store integrity proofs of Rosbag messages, ensuring immutability, simplifying verification, and enhancing data reliability for explanation generation.
2.4 Explainable Autonomous Robots
XARs are designed to provide explanations for their actions and decisions to non-expert users. Such explanations enhance security and trust, helping to prevent failures, errors, conflicts, and undesired interactions while achieving objectives [59, 60]. However, despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical usability and effective interpretability in real environments [61].
Effective explanations in XARs should detail alternatives, the choices made, expectations for each option, decision criteria, and any unexpected events that may have led to changes in the planned behavior [17, 62]. These explanations often rely on abductive reasoning, which bridges the gap between the robot’s internal decision-making process and the user’s understanding [63]. To support this, frameworks have been developed that enable robots to articulate their internal states and decisions, enhancing collaboration and improving communication with humans [64].
Adaptive algorithms and needs-based motivational systems allow robots to provide real-time, context-aware explanations tailored to diverse audiences [65, 66]. Behavior Trees (BTs) hierarchically decompose tasks into goals and subgoals, making complex behaviors easier to understand and adaptable to real-time changes [67]. However, achieving a balance between simplicity and adequacy is essential when making complex information easy to understand. By focusing on the most relevant details and filtering non-essential information, explanations foster trust and collaboration by aligning with human cognitive capacities [68]. For this purpose, our architecture incorporates a log curation module that processes raw robotic data, filtering it to retain only meaningful, high-relevance information, thus delivering clearer and more concise explanations directly tied to events.
In addition to context-aware explanations, the core of explainability in XAR lies in effective communication [63]. Progress in Natural Language Processing (NLP) and LLMs can improve the efficiency of human-robot interactions. This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly improving their usefulness and applicability [1].
2.5 Large Language Models in Robotics
LLMs, enabled by self-attention mechanisms, have advanced NLP, improving task-solving capabilities and enabling more effective human-robot interaction [69, 70, 71]. However, LLMs face challenges such as hallucinations, which arise due to reliance on internal knowledge, and the high update costs associated with large parameter sizes, making traditional pre-training and fine-tuning methods impractical. Retrieval-Augmented Generation (RAG) methods address these issues by enabling LLMs to interact effectively with external sources [72, 73]. This approach is particularly advantageous in robotics, where real-time and accurate contextual understanding is essential for complex decision-making and task execution.
Explainability and transparency have become critical aspects of deploying LLMs in autonomous agents. These models enable robots to translate high-level commands into actionable plans, improving performance in complex environments [74, 75, 76]. When integrated with reward decomposition frameworks, LLMs can generate intuitive explanations, linking actions to object-specific properties, thereby reducing ambiguity [77]. LLMs can also generate natural language descriptions of robot behaviors and logs, highlighting opportunities for improvement in accuracy and completeness [78, 79].
Existing solutions, however, often rely on static models or predefined data, limiting their ability to handle dynamic environments. Our approach addresses these limitations by integrating blockchain-supported accountability with curated logging of relevant robotic data. This framework provides robust context for LLM-generated explanations, minimizing hallucinations and ensuring alignment with the robot’s actions and interactions within its environment.
@@ 2.5 Large Language Models in Robotics
Evaluating LLMs in robotics presents additional challenges, including the absence of suitable metrics, biases in generated text, and the high cost of human evaluations [80]. While human evaluations remain the gold standard for capturing subjective insights [48], they are resource-intensive and lack scalability [81]. LLM-based evaluations offer a scalable alternative, achieving over 80% alignment with human judgments while delivering explainable outputs [82]. A hybrid framework combining human assessments for nuanced insights and LLM evaluations for efficiency and repeatability provides a comprehensive and scalable solution for evaluating LLM outputs.
3 System Architecture

### v4

@@ Abstract
The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns. Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behavior to non-expert users. Such explanations are essential to fostering trust and ensuring safety. Additionally, they contribute to improving communication by bridging the gap between the agent and the user, thereby enhancing the effectiveness of their interactions. This paper presents an accountability and explainability architecture implemented for mobile Robot Operating System-based robots. The proposed solution comprises two main components. First, a black box-like element used to provide accountability, featuring anti-tampering properties achieved through blockchain technology. Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models over the data contained within the previously mentioned black box. This study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities. This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing the challenges inherent in the use of autonomous agents in real-world scenarios.
Index Terms:
Accountability, explainability, autonomous agents, robotics, blockchain, human–robot interaction, large language models
@@
=-21pt
I Introduction
The increasing integration of mobile robots into human environments has raised significant concerns regarding their transparency and trustworthiness. A key challenge in this context is ensuring that robots can provide meaningful justifications for their actions and decisions, particularly in safety-critical scenarios where their behavior may impact human well-being [1]. Transparent and reliable explanations are essential [2], as they not only support effective human-robot interactions but also help build trust, particularly among non-expert users who depend on clear, accessible interpretations of robotic behavior [3, 4]. However, the complexity of robotic decision-making processes and the need for thorough post-event analysis make it particularly difficult to achieve these goals.
To address these challenges, accountability must be ensured using a robust tamper-proof logging system. The most effective evidence for post-event analysis comes from logging and auditing capabilities, which are essential for mitigating technical issues and providing forensic evidence to detect data tampering or unauthorized breaches. These systems must guarantee the integrity, confidentiality, and availability of data while enabling incident detection, compliance, and forensic investigation [5].
However, covering tracks in log systems has become common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions. Therefore, it is necessary to strengthen the system properties to prevent unauthorized modifications [6]. This includes safeguarding the integrity of log data and implementing measures to detect and counteract tampering attempts. Effective accountability solutions must ensure the detection and isolation of faulty behaviors and their origin. Moreover, they should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
Integrating accountability solutions into robotic systems often demands a trade-off with system performance. Specifically, the computational power and networking bandwidth may be adversely affected [8]. The inclusion of anti-tampering techniques, such as full disk and communication encryption, can be restrictive due to resource limitations [9]. Additionally, high-cost tamper-proof storage devices may prove economically unviable, particularly when dealing with continuously generated log data.
@@ I Introduction
To this end, blockchain solutions have emerged as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability. Consensus protocols established by all blockchain network members enable the detection and rejection of any destruction or manipulation of data through illicit channels. The nature of this technology mitigates tampering risk and safeguards data [11].
@@
Nevertheless, while blockchain secures log data, the evidence obtained poses significant challenges for analysis, due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events. These factors may result in limited usefulness and misalignment with the expectations of non-expert users [12]. The raw log messages produced by autonomous agents are typically noisy and semi-structured. In addition, it is necessary to deal with specific features such as large quantities of domain-specific language messages [13]. Addressing this issue requires methods to organize and filter log data, making it accessible and meaningful.
@@
In this context, explaining the reasons behind a robot’s specific and unexpected actions is essential for diagnosing failures and comprehending deviations from the intended goals. Equally important is ensuring that these explanations are understandable to users. Explainable Artificial Intelligence (XAI) is a research area dedicated to enhancing the transparency and interpretability of Artificial Intelligence (AI) systems to ensure their responsible and ethical deployment. As autonomous robots grow in complexity, understanding and predicting their behavior has become increasingly challenging for non-expert users. This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents [14]. XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques such as natural language explanations to improve clarity and user comprehension [15, 16].
@@
To further improve the effectiveness of XAR systems, Large Language Models (LLMs) are used in diverse natural language processing tasks, including command translation and task planning [17]. Our approach integrates LLMs with blockchain-secured log data. Using retrieval-based techniques, this solution generates accurate and context-aware natural language explanations based on verifiable information [18]. This integration ensures that explanations are clear, coherent, and grounded in a traceable and reliable context, effectively addressing user comprehension and data integrity concerns.
Although recent studies have advanced accountability and explainability in autonomous robotic systems, many existing approaches still face important limitations. Several rely on fixed templates or static rule-based methods that lack flexibility to adapt to dynamic and unpredictable environments. Consequently, the explanations produced may be generic, rigid, or poorly aligned with the actual context of the robot’s behavior. Moreover, explainability techniques are often developed separately from accountability mechanisms, leading to explanations not grounded in verifiable data. This disconnect limits their reliability and effectiveness, particularly in safety-critical settings where trust and traceability are essential.
@@
In addition, current evaluation methods for these explanations face significant challenges, including the absence of standardized and domain-specific metrics, biases in the generated text, and the high cost of human evaluations [19]. Even though human evaluations remain the gold standard for capturing subjective insights [20], they are resource-intensive and lack scalability [21]. Alternatively, LLM-based evaluation methods offer a scalable alternative, demonstrating over 80% alignment with human judgments while offering interpretable and explainable assessments [22, 23].
I-A Contributions
This work presents an accountability and explainability architecture designed for Robot Operating System (ROS)-based mobile robots. This solution integrates two main components. The first is a black box-like module that provides accountability through anti-tampering features by distributing integrity proofs from the events recorded by the black box. The second component generates natural language explanations based on black box data, including interactive natural language conversations as a means of providing comprehensive explainability for the recorded events. Both components are connected through an additional module responsible for processing the raw data collected by the first module, eliminating non-essential information, and building connections and relations between data. These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations to document the code and behaviors of the autonomous agent.
To validate our approach, we conducted a comprehensive evaluation to examine both the impact of the architecture on system performance and the quality of the generated explanations. In order to capture detailed insights while maintaining efficiency and reproducibility, we developed a hybrid evaluation methodology that combines human judgments with LLM-based assessments. The evaluation focused on key criteria such as correctness, accuracy, helpfulness, coherence, and user satisfaction. This integrated framework provides a robust, scalable, and reliable means of assessing LLM-generated explanations in robotic systems. Our findings demonstrate that the proposed architecture successfully balances accountability and explainability without compromising robotic performance, thereby offering a practical and scalable solution for deployment in real-world applications.
Therefore, the primary contributions of this work are summarized as follows:
  1. 1.
    A unified architecture that combines blockchain-based tamper-proof logging with LLMs to generate grounded natural language explanations of robot behavior, enabling accountability and transparency.
  2. 2.
    Real-time integrity and verifiability of robotic event data through a blockchain-based black box component, with minimal impact on robot performance, even at high data rates, addressing a critical requirement for practical deployment.
  3. 3.
    A log curation and processing module that filters noisy, semi-structured robotic logs and extracts meaningful context, significantly improving the relevance, structure, and coherence of the generated explanations.
  4. 4.
    An empirical analysis of the trade-offs between system performance and logging granularity, offering practical design insights for developing robust, interpretable, and efficient autonomous agents
  5. 5.
    A hybrid and comprehensive evaluation methodology combining LLM-as-judge and human-based assessments, designed to leverage the strengths of both approaches in validating the quality, accuracy, and user satisfaction of the generated explanations for complex robotic behaviors such as goal cancellations, re-planning, and obstacle avoidance.
I-B Structure of the Paper
The remainder of this paper is organized as follows: Section II provides an overview of previous studies and findings related to the study, focusing on the enhancement of immutable data recorders and explainability solutions for robots. Section III delves into the system architecture, offering a detailed description of the designed approaches. Section IV describes the technical elements of the system designed for the proposed architecture. Section V presents the evaluation process, including the scenarios and assessment variables. Section VI outlines the results obtained, particularly addressing accountability and explainability metrics. Finally, Section VII summarizes the obtained conclusions.
II Related Works
In this section, we explore key works related to various domains relevant to our approach. This includes advances in how blockchain ensures the immutability of logs, research on applying Event Data Recorders (EDRs) in robotics, significant contributions to integrating XAI with blockchain, and XAR. Finally, we examine the evolution of LLMs to generate natural language explanations and their advancements in robotics.
II-A Blockchain for Immutable Logs
Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks. These blocks are linked through the hashing and storage of the previous block content, ensuring the integrity of transactions. Any attempt to alter a transaction in one block requires modification of all subsequent blocks, resulting in a computationally expensive process. As a result, the data stored in the blockchain is immutable [24, 25].
There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, whereas permissioned blockchains limit participation and transaction visibility. Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of block modification [26]. While permissionless blockchains have traditionally been related to scalability and throughput challenges, approaches such as Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible, and scalable consensus protocol [27, 28].
Blockchain technology ensures log integrity and auditability, offering unique advantages over alternatives such as distributed storage systems or InterPlanetary File Systems (IPFSs), which may require central authorities for storage and verification [29]. Blockchain offers consensus and immutability, coupled with the capability of integrating logic mechanisms such as Smart Contracts (SCs). SCs contain functions that either write values onto the blockchain, thus altering its state, or reading data from it [30]. These features make blockchain the preferred choice for applications in which data integrity and trust are required [31].
Different secure logging systems rely on permissioned blockchain frameworks for log immutability [32], introducing accountability even in embedded systems [7]. To achieve a balance between throughput and cost-effectiveness, some works combine both categories [33], developing approaches that include a permissioned blockchain to ensure per-entry immutability and a public non-repudiation solution [34]. Nevertheless, inherent features of permissionless blockchains such as transparency and trust, have boosted the development of secure logging systems [35].
While previous work often records cryptographic hashes on the blockchain to ensure log integrity and employs SCs for data management, our approach goes further by using these verified records as contextual input for generating natural language explanations of autonomous robot behavior. Additionally, our system includes features that enable public accountability and prevent denial of recorded actions. This design provides strong guarantees against data manipulation and ensures that the explanations produced by our system are always based on trustworthy, tamper-proof information. By bridging immutable records with interpretable explanations, our method enhances transparency and clarity in communicating robot actions.
II-B Blockchain for Event Data Recorders in Robotics
EDRs, or black boxes, are essential for capturing and analyzing events in robotic systems. These recorders help comprehend system behavior, detect anomalies, and ensure safety [36, 37]. Therefore, these solutions are required for forensic investigations and ethical assessments of autonomous agents [38]. The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [39, 9]. These solutions typically include hashing, signing, and encryption [40], while ensuring event integrity through anti-tamper mechanisms [41].
The blockchain technology in EDRs ensures the integrity of a transparent and auditable record of events. These features enhance accountability and trust in the recorded data by preventing tampering [42]. Decentralized and distributed storage further eliminates reliance on central authorities and mitigates data loss risks.
Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management. The Black Block Recorder (BBR) system uses permissioned blockchain for immutable logging, combining Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity [43]. Blockchain applications in robot swarms, including secure communication, data logging, and consensus agreements, are discussed in [44]. In [45], a time-segmented consortium blockchain for robotic event registration is proposed, enabling efficient event organization and retrieval within specific time frames. In [46], the authors present a tool to bridge ROS and Ethereum blockchain, evaluating encryption efficiency and stability, transaction response time, and packet loss rate.
Existing blockchain-based logging solutions in robotics usually focus on low-level data storage and network performance metrics, and often lack mechanisms to record integrity proofs of structured and filtered messages at configurable intervals. Furthermore, these solutions generally miss granular recording policies and the use of SCs to improve data verifiability. In contrast, our approach guarantees tamper-proof evidence by periodically anchoring integrity proofs of curated log data on the blockchain, while also generating natural language explanations grounded in verifiable and filtered information. This integration significantly improves both the accountability of robotic systems and the comprehensibility of the explanations provided to users in dynamic contexts.
II-C Explainable Artificial Intelligence and Blockchain
XAI aims to align AI decisions with human cognitive expectations by employing techniques such as contextual relationships, contrastive explanations, counterfactual reasoning, and interpretable models that enhance interpretability and transparency [47, 48]. Insights from social sciences show that people tend to interpret explanations more positively when they reflect human-like reasoning, emphasizing intentions, goals, and contextual factors rather than purely technical or statistical descriptions [20]. By combining XAI with transparent blockchain record-keeping, explanations can be retraced and validated, ensuring both interpretability and accountability in AI-driven systems [49, 50].
The benefits of this integration are remarkable in several domains. In healthcare, XAI and blockchain have been combined for secure, transparent patient data management, logical disease prediction, and immutable medical diagnostic records [51]. Similarly, in legal domain applications, blockchain and XAI support tamper-proof decision auditing, using IPFS and Ethereum for cryptographic proofs that enable data verification and accountability [52, 29]. Privacy-preserving systems with homomorphic encryption and semantic ontologies further ensure data protection, while maintaining explainability [53].
Regarding autonomous agents, such as unmanned aerial vehicles or autonomous vehicles, blockchain and XAI ensure secure and transparent decision-making [54]. This integration enhances collaboration, even in complex and time-critical scenarios [55]. Decentralized edge models for collaborative AI leverage blockchain for secure communication and real-time data sharing, further optimizing resources, and enhancing reliability in these environments [56].
Despite previous advances, providing explanations grounded in robot event data remains challenging due to the complexity, volume, and heterogeneity of the generated information. To address this, our proposal introduces Rosbag2 files as an alternative to IPFS or distributed file systems for storing contextual data in autonomous robotic systems. This approach supports reproducibility and enables a more precise analysis of robotic behaviors [43]. Rosbag files provide a standardized format for recording large volumes of robotic data, facilitating comprehensive interaction capture for improved analysis. To mitigate the inherent indeterminism of LLM algorithms, we propose using a permissionless blockchain to store integrity proofs of Rosbag messages, ensuring data immutability, simplifying verification, and increasing reliability in explanation generation.
II-D Explainable Autonomous Robots
XARs are designed to provide explanations of their actions and decisions to non-expert users. Such explanations enhance security and trust, helping to prevent failures, errors, conflicts, and undesired interactions while achieving objectives [57, 58]. However, despite the growing demand for explainability in autonomous agents, their use lacks practical usability and interpretability in real environments [59].
Effective explanations in XARs should detail the alternatives, choices made, expectations for each option, decision criteria, and any unexpected events that may have led to changes in the planned behavior [16, 60]. These explanations often rely on abductive reasoning, which bridges the gap between the robot’s internal decision-making process and the user’s understanding [61]. To support this, frameworks have been developed that enable robots to articulate their internal states and decisions, enhance collaboration and improve communication with humans [62].
Adaptive algorithms and needs-based motivational systems allow robots to provide real-time, context-aware explanations tailored to diverse audiences [63, 64]. Behavior Trees (BTs) hierarchically decompose tasks into goals and subgoals, making complex behaviors easier to understand and adaptable to real-time changes [65]. Nevertheless, achieving a balance between simplicity and adequacy remains a key challenge when conveying complex or technical information in an understandable way. Explanations that are overly detailed can overwhelm users, while those that are too simplified may omit critical context necessary for constructing accurate and meaningful interpretations [66].
Our architecture addresses these limitations by incorporating a log curation module that automatically processes and filters raw robotic data, retaining only the most relevant information for explanation generation. By emphasizing essential details and excluding non-essential information, our system produces explanations that align with human cognitive capacities, thereby fostering trust and facilitating effective collaboration.
In addition to providing context-aware explanations, effective explainability in XAR fundamentally relies on clear communication [61]. While prior research has acknowledged the potential of Natural Language Processing (NLP) and LLMs to enhance human-robot communication [61, 1], their integration with robot-specific data streams remains limited. This limitation arises because LLMs primarily depend on probabilistic patterns rather than direct access to data from the robot, which can lead to responses that are linguistically plausible yet factually inaccurate. Our approach addresses this challenge by grounding explanations in curated robotic data and leveraging the natural language capabilities of LLMs, thereby producing clear and accessible explanations that enhance transparency and foster trust between users and autonomous systems, including non-expert audiences.
II-E Large Language Models in Robotics
LLMs, enabled by self-attention mechanisms, have advanced NLP, improving task-solving capabilities and enabling more effective human-robot interaction [67, 68]. However, LLMs face challenges such as hallucinations, which arise due to reliance on internal knowledge, and the high update costs associated with large parameter sizes, making traditional pre-training and fine-tuning methods impractical. Retrieval-Augmented Generation (RAG) methods address these issues by enabling LLMs to interact effectively with external sources [69, 70]. This approach is particularly advantageous in robotics, where real-time and accurate contextual understanding is essential for complex decision-making and task execution.
Explainability and transparency have become critical for deploying LLMs in autonomous agents. These models enable robots to translate high-level commands into actionable plans, improving their performance in complex environments [71, 72, 73]. When integrated with reward decomposition frameworks, LLMs can generate intuitive explanations, linking actions to object-specific properties, thereby reducing ambiguity [74]. LLMs can also generate natural language descriptions of robot behaviors and logs, highlighting opportunities for improvement in accuracy and completeness [75, 76].
However, existing solutions often rely on static models or predefined data, which limit their ability to handle dynamic environments. Furthermore, these solutions rarely provide mechanisms for ground explanations in verifiable, tamper-proof evidence of robot behaviors. Our approach addresses these limitations by integrating blockchain-supported accountability with the curated logging of relevant robotic data. This framework provides a robust context for LLM-generated explanations, minimizes hallucinations, and ensures alignment with the robot’s actions and interactions within its environment.
III System Architecture